OpenStudy (anonymous):

find the derivative of tan inverse ( (sqrt(1+x^2)+sqrt(1-x^2) ) / ( (sqrt(1+x^2)-sqrt(1-x^2) )

OpenStudy (anonymous):

multiply by the conjugate

hartnn (hartnn):

i am sure there is some tremendous simplification involved before taking the derivative. try and put x= tan y \(\huge \dfrac{\sqrt{1+x^2}+\sqrt{1-x^2}}{\sqrt{1+x^2}-\sqrt{1-x^2}}=\dfrac{1+\sqrt{\dfrac{1-x^2}{1+x^2}}}{1-\sqrt{\dfrac{1-x^2}{1+x^2}}}\) with x= tan y \(\sqrt{\dfrac{1-x^2}{1+x^2}}=\sqrt{\dfrac{1-\tan^2y}{1+\tan^2y}}=...?\) can you simplify this ?

OpenStudy (anonymous):

no i'm not getting this simplified but long route did give me answer as - x / sqrt(1-x^4)

hartnn (hartnn):

you tried with conjugate right ?

hartnn (hartnn):

actually, x^2 =sin y will be a better substitution

hartnn (hartnn):

\( \large \sqrt{\frac{1 - \sin\theta}{1 + \sin\theta}} = \frac{1 - \tan(\theta/2)}{1 + \tan(\theta/2)} \)

OpenStudy (anonymous):

no i tried directly diff for tan inverse with chain rule

hartnn (hartnn):

with x^2 = sin y \(\huge \dfrac{\sqrt{1+x^2}+\sqrt{1-x^2}}{\sqrt{1+x^2}-\sqrt{1-x^2}}=\dfrac{1+\sqrt{\dfrac{1-x^2}{1+x^2}}}{1-\sqrt{\dfrac{1-x^2}{1+x^2}}} \\ \huge \dfrac{1+\dfrac{1-\tan(y/2)}{1+\tan(y/2)}}{1-\dfrac{1-\tan(y/2)}{1+\tan(y/2)}}=\dfrac{1}{\tan(y/2)} \\ \huge \tan^{-1}(\cot(y/2))=\pi/2-y/2 = \\\huge =\pi/2-\sin^{-1}(x^2)/2\) see, how simplified it became :)

hartnn (hartnn):

and then one step derivative of sin inverse will give you required answer, same as what u got :)

hartnn (hartnn):

got how i simplified? ask if any doubts :)

OpenStudy (anonymous):

u said x^2 = sin y, then how did u get tan y/2

OpenStudy (souvik):

1-siny = sin^y/2+cos^y/2-2siny/2*cosy/2 =(cosy/2-siny/2) ^2 similarly 1+siny = (cosy/2+siny/2)^2

OpenStudy (anonymous):

thanks a lot

OpenStudy (souvik):

you are welcome..